OpenStudy (anonymous):

divide (y^2-3y+2)/(y^2+5y+6) / (y^2+2y-8)/(y^2+y-12) use complete sentences to explain how to simplify this?

OpenStudy (mathmath333):

\[\text{factor all of these, ex.}\\y^2-3y+2\\=y^2-y-2y+2\\=y(y-1)-2(y-1)\\=(y-1)(y-2)\]

OpenStudy (danjs):

I am assuming this is it \[\frac{ y^2 - 3y +2 }{ y^2 + 5y +6 } *\frac{ y^2 + y -12 }{ y^2+2y - 8 }\]

OpenStudy (danjs):

\[\frac{ (y-2)(y-1) }{ (y+3)(y+2) }*\frac{ (y+4)(y-3) }{ (y+4)(y-2) }\]

OpenStudy (danjs):

Cancel out, (y+4) , (y-2) \[\frac{ (y-1)(y-3) }{ (y+3)(y+2) }\]

OpenStudy (danjs):

\[\frac{ y^2 - 4y + 3 }{ y^2 + 5y +6 }\]

OpenStudy (danjs):

1) multiply by the reciprical of the bottom fraction 2)Factor each polynomial 3) cancel common terms, ratios of 1 4) distribute out the remaining terms, if you want to.

OpenStudy (danjs):

This was an old post, figure i would just complete it for you to review later.